# ONG: Orthogonal Natural Gradient Descent

### v1

@@ Abstract
Orthogonal gradient descent has emerged as a powerful method for continual learning tasks. However, its Euclidean projections overlook the underlying information-geometric structure of the space of distributions parametrized by neural networks, which can lead to suboptimal convergence in learning tasks. To counteract this, we combine it with the idea of the natural gradient and present ONG (Orthogonal Natural Gradient Descent). ONG preconditions each new task gradient with an efficient EKFAC approximation of the inverse Fisher information matrix, yielding updates that follow the steepest descent direction under a Riemannian metric. To preserve performance on previously learned tasks, ONG projects these natural gradients onto the orthogonal complement of prior task gradients. We provide a theoretical justification for this procedure, introduce the ONG algorithm, and benchmark its performance on the Permuted and Rotated MNIST datasets. All code for our experiments/reproducibility can be found here at https://github.com/yajatyadav/orthogonal-natural-gradient.
@@ 1 Introduction
Continual learning, the process of training a single model on a sequence of tasks without forgetting previously learned tasks, is a big challenge in deep learning. Naive fine-tuning leads to an issue known as catastrophic forgetting, where earlier tasks’ performance collapses as NNs shift their optimal parameters and weights to accommodate the new task, forgetting previously learned tasks (McCloskey and Cohen, 1989; Goodfellow et al., 2015). In deep learning, a promising method to tackle this problem is orthogonal gradient descent. Orthogonal gradient descent proposes a solution where each new-task gradient is projected into an orthogonal subspace spanned by previous tasks’ gradient steps. By subtracting off this projection before taking the gradient step, we ensure that the model predictions aren’t affected by the trained parameters of previous tasks. Another idea in optimization is the natural gradient. Natural gradients arise when accounting for the underlying geometry of the parameter space in our optimization to find the steepest direction. Natural gradient descent has been found to require far fewer iterations than traditional gradient descent in many applications. It is also invariant to any smooth and invertible reparameterization of the model, as compared to gradient descent, which is parametrization dependent. Natural gradient descent works by modifying the gradient update rule to account for the "information geometry" of the parameter space. In the context of probabilistic models, this involves preconditioning the gradient in the update rule with the inverse Fisher information matrix, as it can be shown that the Riemannian metric structure of this parameter space is described by the Fisher information. This new formulation ensures optimization in the steepest descent direction. Our paper aims to incorporate natural gradients in the OGD algorithm to combine the strengths of these methods and create a robust continual learning algorithm. We present our algorithm, Orthogonal Natural Gradients (ONG), and provide some theoretical guarantees and an efficient implementation. We then test it on standard continual learning benchmarks.
@@ 2 Related Work
Continual learning methods broadly fall into three categories: regularization-based, replay-based, and dynamic architecture methods. Regularization-based methods add task-specific penalties to the loss to preserve old knowledge. Notable works include Elastic Weight Consolidation (Kirkpatrick et al., 2017), which uses the Fisher information at a task’s optimum as a quadratic penalty on parameter changes, and Synaptic Intelligence (SI) (Zenke et al., 2017), which integrates a Hebbian-like importance of each weight over tie. Memory Aware Synapses (MAS) compute parameter importance by measuring the sensitivity of the output function to each weight, and penalizes updates to important parameters (Aljundi et al., 2018). All of these effectively “slow down” learning on critical weights after each task. Recent variants have hybridized and/or extended on these ideas (Schwarz et al., 2018; Loo et al., 2020; Batra and Clark, 2024). Replay-based methods mitigate forgetting by rehearsing data from previous tasks during new-task training. This can be done by either storing real data or generating synthetic examples that approximate prior distributions. Deep Generative Replay (DGR) methods (Shin et al., 2017) train a generative model to mimic old data distributions. During new task training, the generative model samples pseudo-data from prior tasks, allowing the learner to “replay” previous experiences without storing real examples. Experience Replay (ER) (Rolnick et al., 2019) stores a subset of training samples from past tasks in a memory buffer and interleaves them with new data during learning. While effective, its drawbacks lie within concerns about memory usage and privacy. Further improvement include Prioritized Generative Replay (Wang et al., 2025) where random samples are generated condition on “relevance” functions, making samples richer and more informative than DGR. Hybrid approaches have also been proposed, combining replay and regularization (Prabhu et al., 2020; Yan et al., 2021). Finally, dynamic architecture methods seek to address catastrophic forgetting by modifying the model structure over time, often by allocating task-specific sub-modules or expanding capacity to accommodate new tasks. Progressive Neural Networks (PNNs) (Rusu et al., 2022) allocate a new column for each task and connect it previous columns via lateral connections. This guarantees no forgetting but scales poorly with the number of tasks. Dynamically Expandable Networks (DEN) (Yoon et al., 2018) grow the network only when necessary, by splitting and duplicating neurons based on task specific losses, allowing for more compact architecture compared to PNNs. Similarly, PackNet (Mallya and Lazebnik, 2018) prunes a large model after each task and reuses freed weights for new tasks, enabling weight reuse without interference. Other dynamic methods use conditional computation or gating mechanisms to route different tasks through different subsets of the network, as in Context-dependent Gating (Masse et al., 2018), which applies task-specific masks over weights to isolate task representations. A notable continual learning method, one of which forms the foundation upon which our work builds and extends on, is Orthogonal Gradient Descent (OGD), originally proposed by Farajtabar et al. (2019). OGD mitigates forgetting by projecting incoming gradients onto the orthogonal complement of the subspace spanned by previously observed gradients. Follow-up work by Bennani et al. (2020) offers theoretical guarantees for OGD in the neural tangent kernel (NTK) regime, demonstrating that under mild assumptions, OGD can preserve task performance indefinitely.
@@
The natural gradient (Amari, 1998) generalizes standard gradient descent by taking steps that respect the geometry of the parameter space. Instead of updating parameters in the Euclidean space, natural gradients optimize in the space of probability distributions using the Fisher information matrix (FIM) as a Riemannian metric. In probabilistic models and Bayesian neural networks, natural gradient descent yields faster convergence and more stable updates by accounting for parameter sensitivity (Khan and Lin, 2017; Martens, 2020). Methods such as Variational Online Gauss-Newton (VOGN) (Khan and Lin, 2017) and Bayesian Gradient Descent (Liu et al., 2023) utilize natural gradients for continual learning by updating variational parameters using FIM-preconditioned gradients. This is conceptually similar to OGD’s gradient projection, in that both methods transform the gradient based on local curvature or constraints, aiming to minimize forgetting.
Efficient Approximations to Fisher Information
While natural gradient methods are theoretically appealing, computing and inverting the full Fisher information matrix is often intractable for deep networks. To address this, several efficient approximations have been proposed: Diagonal approximations (as in standard EWC) are simple but ignore parameter interactions. Kronecker-Factored Approximate Curvature (K-FAC) (Martens, 2020) factorizes the FIM into smaller Kronecker components, significantly reducing storage and compute requirements. EK-FAC (Eigenvalue-Corrected K-FAC) (George et al., 2018) improves on K-FAC by computing the eigenbasis of each factor and tracking variances in that space. This method provably offers a closer approximation to the true FIM in the Frobenius norm while retaining computational efficiency. Ritter et al. (2018) allow continuous curvature estimation using block-diagonal structures, further aiding in scalability for continual learning.
@@ 3.1 Continual Learning
In continual learning, a model encounters tasks $\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3},$... sequentially, each of which describes their own supervised learning task. Once the data task for $\mathcal{T}_{k}$ has been utilized for training, it is usually discarded (each example is seen once). The challenge of this method lies in updating the model so that it performs well on the current task without sacrificing its ability to solve all previously learned tasks (Nguyen et al., 2017). Formally, we assume that for each task $\mathcal{T}_{k}$, the data $\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{n_{k}}$ are drawn i.i.d. from some distribution $\mathcal{D}_{k}$ (Bennani et al., 2020). We denote the model parameters after $t$ iterations on task $k$ by $f_{k}^{(t)}$, and then use $f^{*}_{k}$ to refer to the final parameters once the model learns the task (Tang et al., 2020).
@@ 3.2 Probabilistic Setup
where $f_{\text{c}}(\boldsymbol{x};\boldsymbol{w})$ is the logit class for c. Our objective is to maximize the log-likelihood (equivalent to minimizing cross-entropy) so the learning problem lives on the manifold of all these predictive distributions. This is what motivates the use of the Fisher information matrix (Amari, 1998):
@@
Note that “Riemannian structure is given to the parameter space of multilayer networks by the Fisher information matrix” as found from Amari (1998). Intuitively, $\boldsymbol{F}(\boldsymbol{w})$ measures the sensitivity of the model’s output distribution to small changes in each parameter direction. Directions that correspond to large Fisher curvature produce large changes in the predictive distribution (which also implies a large KL divergence between the old and new probability distributions after taking a step in parameter space), while the “flat” directions do not affect the model’s behavior much. This geometric structure on our parameter space motivates the natural gradient
$\delta=\boldsymbol{F}^{-1}(\boldsymbol{w})\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})$ (3)
as being the steepest-descent direction of loss function $\mathcal{L}(\boldsymbol{w})$ over our manifold. The preconditioning by the inverse Fisher matrix also provides us with optimization steps that are invariant to model reparameterization, which can be useful in tasks such as classification (Martens, 2020).
@@ 3.3 Orthogonal Gradient Descent
Algorithm 1 Orthogonal Gradients Descent
@@
The above algorithm is the original OGD algorithm  Farajtabar et al. (2019). To justify the use of orthogonal gradient descent, it has been shown that every model update $\tilde{\boldsymbol{g}}$ remains a valid descent direction and that the generalization gap can be bounded just as tightly as it is for SGD (Bennani et al., 2020). The forgetting bound is also tightly controlled by a factor that can limit how much any one past task’s loss can rise. Together, these results show that OGD is not just empirically effective, but also comes with provable convergence, bounded generalization error, and guaranteed forgetting control  Farajtabar et al. (2019). All these aspects make it theoretically sound and appealing for continual learning and our applications with the natural gradient.
@@ 4 Method
We now describe our methodology to combine the core ideas of OGD and natural gradients to create a new continual learning algorithm ONG: Orthogonal Natural Gradient. This algorithm combines the parameter space invariant updates of natural gradients with the interference-free projections of orthogonal gradient descent (OGD). The key idea is partitioned into several steps.
@@ 4.1 Natural Preconditioning
For each new task gradient, $\boldsymbol{g}=\nabla_{\boldsymbol{w}}\mathcal{L}_{T_{k}}$, we edit this gradient to instead be
$\boldsymbol{g}=\boldsymbol{F}^{-1}\ \nabla_{\boldsymbol{w}}\mathcal{L}_{T_{k}}$ (4)
where $\boldsymbol{F}$ is the approximate Fisher information matrix, estimated for the current set of model parameters. This step moves the parameters $\boldsymbol{w}$ along the steepest direction on the manifold of the predictive distributions, which minimizes the KL-divergence from the conditional distribution parameterized by the previous model and ensures reparameterization invariance. The remaining steps from the previous algorithm remain the same. We still maintain a basis $S$ of “sensitive” directions and project $\boldsymbol{g}$ into the orthogonal complement of the span of $S$:
@@
We perform the same parameter update
$\mathbf{w}\;\leftarrow\;\mathbf{w}-\eta\,\tilde{\mathbf{g}}$,
and once task $T_{k}$ converges—just as in the OGD paper—we extract each logit-wise gradient $\nabla f_{k}(\boldsymbol{x};\,\boldsymbol{w})$. Then, as before, we precondition this gradient, project it onto the orthogonal complement of the subspace spanned by all previously added logit-wise gradients, and denote the result by $\boldsymbol{u}$. Finally, we add $\boldsymbol{u}$ to our set $S$, maintaining $S$ as an orthogonal basis of “sensitive directions.” $S$ spans the subspace of all previously seen natural gradients.
@@ 4.2 OGD-Plus (OGD+)
OGD-Plus(OGD+) is an algorithm that enhances vanilla OGD by storing the feature maps w.r.t samples from previous tasks, in addition to feature maps w.r.t samples from current task (which is what OGD stores) (Bennani et al., 2020). These extra samples are saved in a dedicated memory, the sample memory. The authors motivated this change in order to make OGD more robust to the NTK (Neural Tangent Kernel) variation phenomena in the experiments. The changes from OGD to OGD+ can be found in blue in the below algorithm.
@@
8:   $\tilde{\boldsymbol{g}}=\boldsymbol{g}-\sum_{\boldsymbol{v}\in S}\text{proj}_{\boldsymbol{v}}(g)$
@@
13:   $u\leftarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{F}^{-1}}\nabla f_{k}(\boldsymbol{x};\boldsymbol{w})-\sum_{\boldsymbol{v}\in S}\text{proj}_{\boldsymbol{v}}\left(\nabla f_{k}(\boldsymbol{x};\boldsymbol{w})\right)$
14:   $S\leftarrow S\cup\{\boldsymbol{u}\}$
15:  end for
16:  Sample $D\subset T_{k}$
17:  Update $S_{D}\leftarrow S_{D}\cup D$
18:end for

### v2

@@ Abstract
Orthogonal Gradient Descent (OGD) has emerged as a powerful method for continual learning. However, its Euclidean projections do not leverage the underlying information-geometric structure of the problem, which can lead to suboptimal convergence in learning tasks. To address this, we propose incorporating the natural gradient into OGD and present ONG (Orthogonal Natural Gradient Descent). ONG preconditions each new task-specific gradient with an efficient EKFAC approximation of the inverse Fisher information matrix, yielding updates that follow the steepest descent direction under a Riemannian metric. To preserve performance on previously learned tasks, ONG projects these natural gradients onto the orthogonal complement of prior tasks’ gradients. We provide an initial theoretical justification for this procedure, introduce the Orthogonal Natural Gradient Descent (ONG) algorithm, and present preliminary results on the Permuted and Rotated MNIST benchmarks. Our preliminary results, however, indicate that a naive combination of natural gradients and orthogonal projections can have potential issues. This finding motivates continued future work focused on robustly reconciling these geometric perspectives to develop a continual learning method, establishing a more rigorous theoretical foundation with formal convergence guarantees, and extending empirical validation to large-scale continual learning benchmarks. The anonymized version of our code can be found as the zip file here.
@@ 1 Introduction
Continual learning, the process of training a single model on a sequence of tasks without forgetting previously learned tasks, is a major challenge in deep learning. Naive fine-tuning leads to catastrophic forgetting, where earlier tasks’ performance collapses as networks shift their weights to accommodate the new task, forgetting previously learned tasks McCloskey and Cohen (1989); Goodfellow et al. (2015). In deep learning, a promising training-time algorithm to mitigate this issue is Orthogonal Gradient Descent (OGD) Farajtabar et al. (2019). OGD works by projecting each task gradient to an orthogonal subspace spanned by previous tasks’ gradients. By subtracting off this projection before taking the gradient step, we ensure that the model retains performance on previously seen tasks. Another idea in optimization is the natural gradient. Natural gradients represent the steepest descent direction with respect to the underlying geometry of the parameter space. Iteratively following the natural gradient yields the natural gradient descent algorithm Amari (1998). In many applications, Natural Gradient Descent requires far fewer iterations to converge than standard gradient descent. It is also invariant to any smooth and invertible reparameterization of the model, as compared to gradient descent, which is parametrization dependent Martens (2020). Natural Gradient Descent works by modifying the gradient update rule to account for the "information geometry" of the parameter space. In the context of probabilistic models, this involves preconditioning the gradient in the update rule with the inverse Fisher information matrix (since the Riemannian metric structure of the parameter space is described by the Fisher information Amari (1998)). Natural Gradient Descent methods have been used in several applications, such as reinforcement learning Kakade (2001) and variational inference Wu and Gardner (2024).
In this work, we present Orthogonal Natural Gradient Descent (ONG), a novel algorithm for continual learning that integrates natural gradients within the Orthogonal Gradient Descent (OGD) framework. ONG is designed to harness the complementary strengths of these methods: the explicit forgetting prevention of OGD and the efficient, geometrically-informed updates of natural gradient optimization. We formalize the ONG algorithm, establish theoretical guarantees for its performance, and provide an efficient implementation. However, our empirical results on standard continual learning benchmarks suggests that naively combining natural gradients with orthogonal projections does not mitigate catastrophic forgetting, and inspires us to continue future work on rigorously developing and justifying the correct way to perform vector-space projection with natural gradients. We also aim to extend our evaluation criteria to more challenging, realistic continual learning settings.
@@ 2 Related Work
Continual learning methods broadly fall into three categories: regularization-based, replay-based, and dynamic architecture methods. Regularization-based methods add task-specific penalties to the loss to preserve old knowledge. Notable works include Elastic Weight Consolidation Kirkpatrick et al. (2017), which uses the Fisher information at a task’s optimum as a quadratic penalty on parameter changes, and Synaptic Intelligence (SI) Zenke et al. (2017), which integrates a Hebbian-like importance of each weight over time. Replay-based methods mitigate forgetting by rehearsing data from previous tasks during new-task training. This can be done by either storing real data or generating synthetic examples that approximate prior distributions. For example, deep Generative Replay (DGR) methods Shin et al. (2017) train a generative model to mimic old data distributions. During new task training, the generative model samples pseudo-data from prior tasks, allowing the learner to “replay” previous experiences without storing real examples. Finally, dynamic architecture methods seek to address catastrophic forgetting by modifying the model structure over time, often by allocating task-specific sub-modules or expanding capacity to accommodate new tasks. For instance, Progressive Neural Networks (PNNs) Rusu et al. (2022) allocate a new column for each task and connect it to previous columns via lateral connections. This guarantees no forgetting but scales poorly with the number of tasks.
A notable continual learning method, which inspires the foundation of our work, is Orthogonal Gradient Descent (OGD), originally proposed by Farajtabar et al. (2019). OGD mitigates forgetting by projecting incoming task gradients onto the orthogonal complement of the subspace spanned by previous tasks’ observed gradients. Follow-up work by Bennani et al. Bennani et al. (2020) offers theoretical guarantees for OGD in the neural tangent kernel (NTK) regime, demonstrating that under mild assumptions, OGD can preserve task performance indefinitely.
@@
Natural Gradient Descent Amari (1998) generalizes vanilla gradient descent by taking steps that respect the geometry of the parameter space. Instead of updating in the Euclidean parameter-space, natural gradients perform optimization in the space of probability distributions, where the Fisher information matrix (FIM) serves as a Riemannian metric. In probabilistic models and Bayesian neural networks, natural gradient descent yields faster convergence and more stable updates by accounting for parameter sensitivity Khan and Lin (2017); Martens (2020).
@@ 3.1 Continual Learning
In continual learning, a model encounters tasks $\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3},$... sequentially, each of which is its own supervised learning problem. Once the data for task $\mathcal{T}_{k}$ has been utilized for training, it is usually discarded (each example is seen once). The challenge lies in updating the model so that it performs well on the current task without sacrificing its ability to solve all previously learned tasks Nguyen et al. (2017). Formally, we assume that for each task $\mathcal{T}_{k}$, the data $\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{n_{k}}$ are drawn i.i.d. from some distribution $\mathcal{D}_{k}$. We denote the model parameters after $t$ iterations on task $k$ by $f_{k}^{(t)}$, and then use $f^{*}_{k}$ to refer to the final parameters of the trained model on all the tasks.
@@ 3.2 Probabilistic Setup
where $f_{\text{c}}(\boldsymbol{x};\boldsymbol{w})$ is the logit for class c. Our objective is to maximize the log-likelihood (equivalent to minimizing cross-entropy) so the learning problem lives on the manifold of all these predictive distributions. This is what motivates the use of the Fisher information matrix Amari (1998):
@@
Amari Amari (1998) demonstrates that a Riemannian structure is given to the parameter space of multilayer networks by the Fisher information matrix. Intuitively, $\boldsymbol{F}(\boldsymbol{w})$ measures the sensitivity of the model’s output distribution to small changes in each parameter direction. Directions that correspond to large Fisher curvature produce large changes in the predictive distribution (which also implies a large KL divergence between the old and new probability distributions), while the “flat” directions do not affect the final distribution much. This geometric structure on our parameter space motivates the natural gradient
$\delta(\boldsymbol{w})=\boldsymbol{F}^{-1}(\boldsymbol{w})\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})$ (3)
as being the steepest-descent direction of loss function $\mathcal{L}(\boldsymbol{w})$ while also constraining the KL-divergence between output distributions. The preconditioning by the inverse Fisher matrix provides us with optimization steps that are invariant to model reparameterization, which can be useful in tasks such as classification Martens (2020).
@@ 3.3 Orthogonal Gradient Descent
Algorithm 1 Orthogonal Gradient Descent
@@
Algorithm 1 (shown above) is the original OGD algorithm Farajtabar et al. (2019). To justify the use of orthogonal gradient descent, it has been shown that every model update $\tilde{\boldsymbol{g}}$ remains a valid descent direction and that the generalization gap can be bounded just as tightly as it is for SGD Bennani et al. (2020). The forgetting bound is also tightly controlled by a factor that can limit how much any one past task’s loss can rise. Together, these results show that OGD is not just empirically effective, but also comes with provable convergence, bounded generalization error, and guaranteed forgetting control Farajtabar et al. (2019). All these aspects make it theoretically sound and appealing for continual learning and our applications with the natural gradient.
@@ 4 Method
We now describe our current progress to combine the core ideas of OGD and natural gradients to form a new continual learning algorithm ONG: Orthogonal Natural Gradient Descent. This algorithm aims to combine the parameter space invariant updates of natural gradients with the interference-free projections of orthogonal gradient descent (OGD). The key idea is partitioned into several steps.
@@ 4.1 Natural Preconditioning
For each new task gradient, $\boldsymbol{g}=\nabla_{\boldsymbol{w}}\mathcal{L}_{T_{k}}$, we modify this gradient to instead be $\boldsymbol{g}=\boldsymbol{F}^{-1}\ \nabla_{\boldsymbol{w}}\mathcal{L}_{T_{k}}$ where $\boldsymbol{F}$ is an approximation of the Fisher information matrix, estimated for the current set of model parameters. This step moves the parameters $\boldsymbol{w}$ along the steepest direction on the manifold of the predictive distributions, which minimizes the KL-divergence from the conditional distribution parameterized by the previous model and ensures reparameterization invariance. The remaining steps from the previous algorithm remain the same. We still maintain a basis $S$ of “sensitive” directions and project $\boldsymbol{g}$ into the orthogonal complement of the span of $S$:
@@
We perform the same parameter update $\boldsymbol{w}\;\leftarrow\;\boldsymbol{w}-\eta\,\tilde{\boldsymbol{g}}$, and, analogous to OGD, once task $T_{k}$ converges, we extract each logit-wise gradient $\nabla f_{k}(\boldsymbol{x};\,\boldsymbol{w})$, precondition this gradient, project it onto the orthogonal complement of the subspace spanned by all previously added logit-wise gradients, and denote the result by $\boldsymbol{u}$. We add $\boldsymbol{u}$ to our set $S$, maintaining $S$ as an orthogonal basis of “sensitive directions.” Thus, the set $S$ thus spans the subspace of all previously seen natural gradients.
@@ 4.2 OGD-Plus (OGD+)
OGD-Plus (OGD+) is an algorithm that enhances vanilla OGD by storing the feature maps with respect to samples from previous tasks, as well as the feature maps with respect to samples from the current task Bennani et al. (2020). These extra samples are saved in a dedicated memory, the sample memory. The authors motivated this change in order to make OGD more robust to the NTK (Neural Tangent Kernel) variation phenomena in their experiments. The changes from OGD to OGD+ can be found in blue in the below algorithm.
@@
8:   $\tilde{\boldsymbol{g}}=\boldsymbol{g}-\sum_{\boldsymbol{v}\in S}\text{proj}_{\boldsymbol{v}}(\boldsymbol{g})$
@@
13:   $\boldsymbol{g}\leftarrow\boldsymbol{F}^{-1}\nabla f_{k}(\boldsymbol{x};\boldsymbol{w})$
14:   $\boldsymbol{u}\leftarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{g}}-\sum_{\boldsymbol{v}\in S}\mathrm{proj}_{\boldsymbol{v}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{g}}\right)$
15:   $S\leftarrow S\cup\{\boldsymbol{u}\}$
16:  end for
17:  Sample $D\subset T_{k}$
18:  Update $S_{D}\leftarrow S_{D}\cup D$
19:end for